Father is 5 times as old as his son.In 18 years the father will twice as old as his son.How is the father at present?
Found 2 solutions by fractalier, tiffany222:
Answer by fractalier(6550)

Let their ages NOW be f and s. So we have
f = 5s and in 18 years
f+18 = 2(s+18)
f+18 = 2s + 36
f = 2s + 18
Now substitute the first into this second one and get
5s = 2s + 18
3s = 18
s = 6 years old
f is then 30 years old

Answer by tiffany222(56)

Let father's age be x
Let son's age be y
x = 5y
x + 18 = 2(y +18)
5y + 18 = 2y +36
3y = 18
y = 6
x = 5y = 5(6) = 30
Final Answer:
Father's Age: 30
Son's Age: 6